OpenStudy (anonymous):

Explain, in complete sentences, how you would factor 180x4 – 125. Provide the factors.

OpenStudy (anonymous):

First, I would take out the GCF of 5 ...

OpenStudy (hoblos):

=5(36x4 - 25) = 5(6x2 - 5)(6x2 +5)